.
........................................................................................an inlet pipe can filla tank in 1 hour, a drain pipe can empty it in 3/2 hours. The inlet pipe was openend at 6:00 a.m. to fill the tank. At 6:20a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling
: .
........................................................................................an inlet pipe can filla tank in 1 hour, a drain pipe can empty it in 3/2 hours. The inlet pipe was openend at 6:00 a.m. to fill the tank. At 6:20a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling

Answer by stanbon(18980)

an inlet pipe can fill a tank in 1 hour, a drain pipe can empty it in 3/2 hours. The inlet pipe was opened at 6:00 a.m. to fill the tank. At 6:20a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling
------------------
Filling pipe rate = 1 job/hour
------------------
Draining pipe rate = 3/2 job/hour
------------------
Together rate: 1-3/2 = -(1/2) job/hour
--------------------
EQUATION:
Both were operating from 6:00am till 6:20am or 1/3 hour.
In that time the tank would stay empty because the drain
pipe is removing water as fast as the inlet pipe puts it
in.
----------------
The inlet pipe will then take 1 hour to fill the tank.
It should be full at 7:20 am.
=====================
Cheers,
Stan H.
Question 154584: .
........................................................................................an inlet pipe can filla tank in 1 hour, a drain pipe can empty it in 3/2 hours. The inlet pipe was openend at 6:00 a.m. to fill the tank. At 6:20a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling
: .
........................................................................................an inlet pipe can filla tank in 1 hour, a drain pipe can empty it in 3/2 hours. The inlet pipe was openend at 6:00 a.m. to fill the tank. At 6:20a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling

Answer by ankor@dixie-net.com(4531)

An inlet pipe can fill a tank in 1 hour; a drain pipe can empty it in 3/2 hours. The inlet pipe was opened at 6:00 a.m. to fill the tank. At 6:20 a.m. someone noticed that the drain has been left open and so he closes it immediately. How long will it take the first pipe to finish filling.
:
Let's deal in minutes here:
Inlet time = 60 min
Drain time = 90 min
:
let t = time (in minutes) required to finish filling the tank.
:
Let the full tank = 1:
:
Total on time for the inlet pipe = (t+20)
On time for the drain = 20 min
:
((t+20))/60 - 20/90 = 1
Multiply equation by 180 to get rid of the denominators, results:
3(t+20) - 2(20) = 180
:
3t + 60 - 40 = 180
3t + 20 = 180
3t = 180 - 20
3t = 160
t = 160/3
t = 53.333 min or 53 min 20 sec to finish filling the tank
:
:
Check with calc (inlet time: 53.33 + 20 = 73.33)
(73.33/60) - (20/90) = 1